find the lag: y= 4cos(3 πx+5)

5/3 to the right
3 to the left
5 to the right
5 / 3 π to the left
Answer #1

standard form

y = Acos(B(x-C))+D where C is lag

so,

y = 4cos(3\pix+5)

rewrite it

y = 4cos(3\pi(x +5/3\pi))+0

on comparing with standard form

lag = 5/3\pi, left ( answer)